OpenStudy (anonymous):

Hi can someone help me please? Show that for every non-zero integer n , the number n^2+3n+2 is not prime

OpenStudy (anonymous):

@liltish007

OpenStudy (ikram002p):

check tow cases 1_ if n is odd n=2k+1 (2k+1)^2+3(2k+1)+2 =4k^2+4k+1+6k+3+2 =4k^2+10k+6=2(2k^2+5k+3) is even so not prime 2_if n is even n=2k (2k)^2+3(2k)+2 =4k^2+6k+2=2(2k^2+3k+1) is even not prime

OpenStudy (anonymous):

The usage of these 2 cases can be used to prove any non prime numbers?

OpenStudy (anonymous):

And for curiosities sake, to determine a prime number in for example prove that f(n)= n-1 is a prime number Can i still use those cases ?

OpenStudy (ikram002p):

hmmm not all its depand but i guess it will help u a lot its kind of number theory .. so u can prove if a fun is always not prime but u cant prove if its always prime f(n)=n-1 is not always prime for example n=9 f(9)=9-1=8 not prime or n=4 f(4)=4-1=3 prime

OpenStudy (anonymous):

I think I get it.. I'll go practice some more .. Thanks for the help you really saved my life :)

OpenStudy (ikram002p):

np .. ur wlc lol !